In 2016, Natural Selection, a nationwide computer dating service, had $538 million of assets and $219 million of liabilities.In 2016, Natural Selection, a nationwide computer dating service, had $538 million of assets and $219 million of liabilities. Earnings before interest and taxes were $139 million, interest expense was $31 million, the tax rate was 40 percent, principal repayment requirements were $25.9 million, and annual dividends were 35 cents per share on 22 million shares outstanding. a. Calculate the following for Natural Selection: (Round your answers to 2 decimal places.) b. What percentage decline in earnings before interest and taxes could Natural Selection have sustained before failing to cover: (Round your answers to 1 decimal place.)
Answer #1

a) calculation of liabilities to equity ratio

Liabilities = $219 million

Equity = assets - liabilities

Equity = $538 - $219 = $319 million

Liabilities to equity ratio = liabilities÷ equity

= $219 million ÷ $319 million = 0.687

Times interest earned ratio = EBIT ÷ INTEREST EXPENSE

= $139 million ÷ $31 million

= 4.48

Times burden covered = EBIT÷ (Interest expense + principal payment × (1- tax rate))

= $139 ÷ ( $31 + $25.9÷(1-40%))

= 1.874

b.

Interest payment requirements%

EBIT needs to declined from $139 million to below $31millionto fail to cover the interest expense

= (139-31)÷ 139 ×100 = 77.69% decline in EBIT

Principal and interest requirement%

EBIT needs to to be declined from$139 million to below$ 74.2 million to fail to cover principal and interest

= ($139-$74.2)÷ 139 ×100 = 46.62%

Principal, interest and common dividend payment%

EBIT needs to be declined from$139 million to below [31+25.9/(1-40%)+(0.35×22)/(1-40%)] =$87 to fail to cover principal, interest and common dividend payments

=($139-$87)÷139×100 = 37.4%